Cyril is two years older than Alyssa. In 7 years, the ratio of their ages is 6:7. Find their ages in 5 years.
Answer by harpazo(655)

c = a + 2
(a+7)/(c+7) = 6/7
7(a+7) = 6(c+7)
7a + 49 = 6c + 42
c = 7
c = a+2
a=5.
(c+5) = 12
(a+5) = 10